# Explainable Graph Neural Networks Under Fire

### v1

@@ Abstract
Predictions made by graph neural networks (GNNs) usually lack interpretability due to their complex computational behavior and the abstract nature of graphs. In an attempt to tackle this, many GNN explanation methods have emerged. Their goal is to explain a model’s predictions and thereby obtain trust when GNN models are deployed in decision critical applications. Most GNN explanation methods work in a post-hoc manner and provide explanations in the form of a small subset of important edges and/or nodes. In this paper we demonstrate that these explanations can unfortunately not be trusted, as common GNN explanation methods turn out to be highly susceptible to adversarial perturbations. That is, even small perturbations of the original graph structure that preserve the model’s predictions may yield drastically different explanations. This calls into question the trustworthiness and practical utility of post-hoc explanation methods for GNNs. To be able to attack GNN explanation models, we devise a novel attack method dubbed GXAttack, the first optimization-based adversarial attack method for post-hoc GNN explanations under such settings. Due to the devastating effectiveness of our attack, we call for an adversarial evaluation of future GNN explainers to demonstrate their robustness.
1 Introduction
@@
Graph Neural Networks (GNNs) have achieved promising results in various learning tasks, including representation learning, node classification, graph classification, link prediction, and anomaly detection (Zhou et al., 2020). However, intricate data representations and non-linear transformations render their interpretation and thus understanding their predictions non-trivial (Amara et al., 2022).
@@
This has led to the introduction of GNN explanation methods, which can improve a model’s transparency and thus increase trust in a GNN model, especially when it is deployed in a decision-critical application (e.g., where fairness, privacy, or safety are important) (Ying et al., 2019). Moreover, they can help identify the scenarios where a GNN model may fail (Vu and Thai, 2020). Explainability of a GNN model can be obtained by designing a self-explainable GNN model, which usually employs a simple model architecture and few parameters, leading to explainability at the cost of suboptimal prediction accuracy; or by extracting post-hoc explanations that do not influence the internal workings of a GNN model and thus maintains its high prediction accuracy. In this paper we only consider post-hoc GNN explanation methods, as many such methods have merged over the past decade, including but not limited to Ying et al. (2019); Pope et al. (2019); Yuan et al. (2020); Luo et al. (2020); Vu and Thai (2020); Bajaj et al. (2021); Yuan et al. (2021); Lucic et al. (2022); Huang et al. (2022).
@@
It has been shown that traditional deep neural networks (DNNs) are vulnerable to adversarial attacks (Szegedy et al., 2013; Chakraborty et al., 2018), where an imperceptible well-designed perturbation to input data can lead to substantially different prediction results. Recently, it was demonstrated that the explanations of predictions made by these DNNs are also fragile (Baniecki and Biecek, 2024). That is, an unnoticeable, well-designed perturbation to input data can result in completely different explanations while the model’s predictions remain unchanged. However, the vulnerability of explanation methods has been primarily studied for DNN models designed for image and text data. In contrast, the vulnerability of GNN explanation methods to adversarial attacks has received very limited attention (Fang et al., 2024; Wang et al., 2024), although a plethora of GNN explanation methods have recently been proposed (Yuan et al., 2022; Li et al., 2022; Kakkad et al., 2023).
@@
Importantly, existing studies have only investigated the (in)stability of GNN explanations under relatively “mild” settings: they only consider random perturbations and/or prediction-altering perturbations. Specifically, Agarwal et al. (2022) theoretically analyzed the stability of three GNN explainers under random perturbations. Further, Wang et al. (2024) considered two scenarios: 1) graphs whose predictions are not changed after randomly flipping $20\%$ edges; and 2) severe perturbations using an off-the-shelf attack algorithm to perturb at most $10\%$ edges to change the GNN predictions (as well as the explanations). Besides, Fan et al. (2023) proposed an adversarial attack to simultaneously change the GNN predictions and their explanations. By altering predictions and explanations, however, the graph can become inherently different. Fang et al. (2024) explored evaluation metrics for GNN explanations from the perspective of adversarial robustness and resistance to out-of-distribution input, where they defined “adversarial robustness" as the difficulty of reversing a GNN prediction by perturbing the complementary of the explanatory subgraph.
@@
Figure 1: When using post-hoc GNN explainers, the explanatory subgraph on a graph with “prediction-preserving" perturbations (right) can strongly differ from that on the original graph (left).
@@
Different from previous studies, we aim to perform perturbations that are both carefully crafted (i.e., optimized rather than random) and imperceptible (i.e., such that the GNN predictions remain unchanged but their explanations differ substantially). As shown in Figure 1, we empirically found that small prediction-preserving perturbations can result in largely different explanations generated by post-hoc GNN explainers. To our knowledge, we are the first to formally formulate and systematically investigate this problem. Specifically, we devise GXAttack, the first optimization-based adversarial attack on post-hoc GNN explanations under this setting. We employ a widely used GNN explainer, PGExplainer (Luo et al., 2020), as example target when designing our attack algorithm. Results on various datasets demonstrate the effectiveness of our approach. Moreover, our experiments show that other widely used GNN explanation methods, such as GradCAM (Pope et al., 2019), GNNExplainer (Ying et al., 2019), and SubgraphX (Yuan et al., 2021), are also fragile under the attacks optimized for PGExplainer. Due to the devastating effectiveness of our attack, we call for an adversarial evaluation of future GNN explainers to demonstrate their robustness.
2 Preliminaries
@@
2.1 Node classification with GNNs
We consider the task of node classification. Given an input graph $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathbf{X}\}$ (or $\mathcal{G}=(\mathbf{A},\mathbf{X})$), a prediction model $f_{\theta}(\cdot)$ assigns a label $c$ from a pre-defined set $\mathcal{C}$ to each node $v\in\mathcal{V}$. In other words, we have $f_{\theta}(\mathcal{G}(v))=c$, where $\mathcal{G}(v)$ denotes the computation graph of node $v$. Typically, $f_{\theta}(\cdot)$ consists of 1) a GNN model used to learn node representations, and 2) a multi-layer perceptron (MLP) classifier that assigns labels. More specifically, given a set of training nodes and their labels $\{(v_{1},y_{1}),...,(v_{n},y_{n})\}$ from graph $\mathcal{G}$, the objective function of a GNN classifier can be defined as
@@ 2.1 Node classification with GNNs
2.2 GNN Explainer
We consider post-hoc GNN explanation methods that provide instance-level explanations: given a graph $\mathcal{G}$, a target node $v$, and prediction $f_{\theta}(\mathcal{G}(v))$, the explainer $g_{\lambda}(v,f(\mathcal{G}(v)))$ aims to ‘explain’ this particular prediction. Following common explanation methods, we assume that explanations are given in the form of an explanatory subgraph, namely $\mathcal{G}_{s}(v)=g(v,f(\mathcal{G}(v)))$, consisting of a small set of important edges and their associated nodes, where the explanatory subgraph is usually connected and edge importance is measured by a so-called edge importance score.
In this paper, we take PGExplainer (Luo et al., 2020) as an example to attack for three reasons. First, Faber et al. (2021) pointed out that GNN explanations providing edge importance are preferable to explanations providing node importance, due to the fact that edges have more fine-grained information than nodes. Second, Agarwal et al. (2023) empirically showed that PGExplainer generates the least unstable explanations, exhibiting a $35.35\%$ reduction in explanation instability relative to the average instability of other GNN explainers. Third, the objective function of PGExplainer is differentiable.
For completeness, we briefly revisit the goal of PGExplainer (see Appendix A for more details). For brevity, we denote $\mathcal{G}(v)$ as $\mathcal{G}$ (and $\mathcal{G}_{s}(v)$ as $\mathcal{G}_{s}$ analogously) throughout the remainder of the paper.
@@ 2.2 GNN Explainer
3 GXAttack: GNN Explanation Adversarial Attacks
Given an input $\mathcal{G}$, a GNN classifier $f_{\theta}(\cdot)$, and a GNN explainer $g_{\lambda}(\cdot)$, the original prediction for the target node $v$ is $f_{\theta}(\mathcal{G}(v))$ and its corresponding original explanatory subgraph is given by $\mathcal{G}_{s}(v)=g_{\lambda}(f_{\theta}(v,\mathcal{G}(v)))$. From an attacker’s perspective, a desirable manipulated graph $\mathcal{G}_{adv}:=\hat{\mathcal{G}}=\phi({\mathcal{G}})$ for attacking node $v$ should have the following properties, where $\phi(\cdot)$ is an attack function that perturbs its input, i.e., it removes and/or adds a small set of edges:
@@ 3 GXAttack: GNN Explanation Adversarial Attacks
    The explanatory subgraph for node $v$ on original graph $\mathcal{G}_{s}(v)=g_{\lambda}(f_{\theta}(v,\mathcal{G}(v)))$ and that on perturbed graph $\hat{\mathcal{G}}_{s}(v)=g_{\lambda}(f_{\theta}(\hat{\mathcal{G}}(v)))$ are substantially different: $\mathcal{G}_{s}(v)\neq\hat{\mathcal{G}}_{s}(v)$. In particular, we aim to make $\hat{\mathbf{A}}_{s}$ strongly different from $\mathbf{A}_{s}$.
3.1 Attack Objectives
@@
$\mathcal{L}:=\text{MI}(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\hat{\mathbf{A}},\mathbf{X}))-\beta*\text{MI}(g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X})),g_{\lambda}(f_{\theta}(\hat{\mathbf{A}},\mathbf{X}))).$ (6)
@@
Final Objective. We directly model the edge perturbation matrix $\mathbf{P}\in\{0,1\}^{n\times n}$, where we flip $\mathbf{A}_{ij}$ (namely $0\to 1$ or $1\to 0$) if $\mathbf{P}_{ij}=1$, and keep $\mathbf{A}_{ij}$ intact otherwise. On this basis, we have
$\underset{\mathbf{P}\in\{0,1\}^{n\times n}\text{~{}s.t.~{}}\sum\mathbf{P}_{ij}<B}{\max}~{}\text{MI}(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\mathbf{A}\oplus\mathbf{P},\mathbf{X}))-\beta*\text{MI}(g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X})),g_{\lambda}(f_{\theta}(\mathbf{A}\oplus\mathbf{P},\mathbf{X}))),$ (7)
@@
3.2 Relaxations and Loss Definitions

### v2

@@ Abstract
Predictions made by graph neural networks (GNNs) usually lack interpretability due to their complex computational behavior and the abstract nature of graphs. In an attempt to tackle this, many GNN explanation methods have emerged. Their goal is to explain a model’s predictions and thereby obtain trust when GNN models are deployed in decision critical applications. Most GNN explanation methods work in a post-hoc manner and provide explanations in the form of a small subset of important edges and/or nodes. In this paper we demonstrate that these explanations can unfortunately not be trusted, as common GNN explanation methods turn out to be highly susceptible to adversarial perturbations. That is, even small perturbations of the original graph structure that preserve the model’s predictions may yield drastically different explanations. This calls into question the trustworthiness and practical utility of post-hoc explanation methods for GNNs. To be able to attack GNN explanation models, we devise a novel attack method dubbed GXAttack, the first optimization-based adversarial white-box attack method for post-hoc GNN explanations under such settings. Due to the devastating effectiveness of our attack, we call for an adversarial evaluation of future GNN explainers to demonstrate their robustness. For reproducibility, our code is available via GitHub.
Index Terms:
Graph Neural Networks, Robustness, Explainable Machine Learning
I Introduction
@@
Graph Neural Networks (GNNs) [1] have achieved promising results in various learning tasks, including representation learning [2, 3], node classification [4, 5], link prediction [6, 7], and anomaly detection [8, 9]. However, intricate data representations and non-linear transformations render their interpretation and thus understanding their predictions non-trivial [10].
@@
This has led to the introduction of GNN explanation methods, which can improve a model’s transparency and thus increase trust in a GNN model, especially when it is deployed in a decision-critical application (e.g., where fairness, privacy, or safety are important) [11]. Moreover, they can help identify the scenarios where a GNN model may fail [12]. Explainability of a GNN model can be obtained by designing a self-explainable GNN model, which usually employs a simple model architecture and few parameters, leading to explainability at the cost of suboptimal prediction accuracy; or by extracting post-hoc explanations that do not influence the internal workings of a GNN model and thus maintains its high prediction accuracy. In this paper we only consider post-hoc GNN explanation methods, as many such methods have merged over the past decade, including but not limited to [11, 13, 14, 15, 12, 16, 17, 18, 19].
@@
It has been shown that traditional deep neural networks (DNNs) are vulnerable to adversarial attacks [20, 21], where an imperceptible well-designed perturbation to input data can lead to substantially different prediction results. Recently, it was demonstrated that the explanations of predictions made by these DNNs are also fragile [22]. That is, an unnoticeable, well-designed perturbation to input data can result in completely different explanations while the model’s predictions remain unchanged. However, the vulnerability of explanation methods has been primarily studied for DNN models designed for image and text data. In contrast, the vulnerability of GNN explanation methods to adversarial attacks has received very limited attention [23, 24], although a plethora of GNN explanation methods have recently been proposed [25, 26, 27].
@@
Importantly, existing studies have only investigated the (in)stability of GNN explanations under relatively “mild” settings: they only consider random perturbations and/or prediction-altering perturbations. Specifically, [28] theoretically analyzed the stability of three GNN explainers under random perturbations. Further, [24] considered two scenarios: 1) graphs whose predictions are not changed after randomly flipping $20\%$ edges; and 2) severe perturbations using an off-the-shelf attack algorithm to perturb at most $10\%$ edges to change the GNN predictions (as well as the explanations). Besides, [29] proposed an adversarial attack to simultaneously change the GNN predictions and their explanations. By altering predictions and explanations, however, the graph can become inherently different. [23] explored evaluation metrics for GNN explanations from the perspective of adversarial robustness and resistance to out-of-distribution input, where they defined “adversarial robustness” as the difficulty of reversing a GNN prediction by perturbing the complementary of the explanatory subgraph.
@@
Figure 1: When using post-hoc GNN explainers, the explanatory subgraph on a graph with “prediction-preserving” perturbations (right) can strongly differ from that on the original graph (left).
@@
Different from previous studies, we aim to perform perturbations that are both carefully crafted (i.e., optimized rather than random) and imperceptible (i.e., such that the GNN predictions remain unchanged but their explanations differ substantially). As shown in Figure 1, we empirically found that small prediction-preserving perturbations can result in largely different explanations generated by post-hoc GNN explainers. To our knowledge, we are the first to formally formulate and systematically investigate this problem under a white-box setting. Specifically, we devise GXAttack, the first optimization-based adversarial white-box attack on post-hoc GNN explanations. We employ a widely used GNN explainer, PGExplainer [15], as example target when designing our attack algorithm. Results on various datasets demonstrate the effectiveness of our approach. Moreover, our experiments show that other widely used GNN explanation methods, such as GradCAM [13], GNNExplainer [11], and SubgraphX [17], are also fragile under the attacks optimized for PGExplainer. Due to the devastating effectiveness of our attack, we call for an adversarial evaluation of future GNN explainers to demonstrate their robustness.
We first summarize related work in Section II. Following this, we provide necessary background in Section III. Then, we formalize the GNN explanation attack problem and introduce our novel attack algorithm GXAttack in Section IV. Next, we describe the experimental setups in Section V. Experiment results and corresponding analyses are given in Section VI. Finally, we conclude the paper in Section VII.
II Related Work
II-A GNN Explanations
[25, 26, 27] perform comprehensive surveys and propose excellent taxonomies of GNN explanation methods. Specifically, GNN explanations methods mainly include: 1) Gradients-based methods: Grad [30], GradCAM [13], GuidedBP [31], Integrated Gradients (IG) [32]; 2) Decomposition-based methods: GraphLRP [33], GNN-LRP [34]; 3) Surrogate-based methods: GraphLIME [19], PGM-Explainer [12]; 4) Generation-based methods: XGNN [14]; 5) Perturbation-based methods: GNNExplainer [11], PGExplainer [15], GraphMask [35], SubgraphX [17]; and 6) CF-based methods: CF-GNNExplainer [18], RCExplainer [16].
II-B Attacks and Defenses of Traditional XAI Methods
It has been shown that traditional XAI methods (namely those designed to explain deep neural networks for image and text data) are susceptible to malicious perturbations [36, 37, 38]. Please refer to [22] for a comprehensive survey.
II-C Robustness/Stability of GNN Explanations
[25, 28] point out that a reliable GNN explainer should exhibit stability, namely minor perturbations that do not impact the model predictions should not largely change the explanations. Particularly, [28] theoretically analyze the stability of three GNN explainers, including vanilla Grad, GraphMask and GraphLIME, for which they derive upper bounds on the instability of explanations. Importantluy, [16, 23, 24, 29] are closely related but are different from our work and they are detailed as follows.
OAR. [23] present a novel metric for evaluating GNN explanations, termed Out-Of-Distribution-resistant Adversarial Robustness (OAR). OAR aims to solve the limitations of current removal- and generative-based evaluations. More concretely, they evaluate post-hoc explanation subgraphs by computing their robustness under attack, which consists of three steps: 1) they formulate the adversarial robustness tailored for GNN explanations problem, which is the minimum adversarial perturbation on the structure of complementary subgraph; 2) they introduce a tractable and easy-to-implement objective of adversarial robustness for GNN explanations; after relaxation, the objective becomes a constrained graph generation problem (namely random perturbations) rather than an adversarial attack problem; 3) they present an Out-Of-Distribution (OOD) reweighting block that aims to confine the evaluation on original data distribution. This work is an initial attempt to explore evaluation metrics for GNN explanations from the perspectives of adversarial robustness and resistance to OOD. In contrast, we focus on optimization-based adversarial attack on post-hoc GNN explanations rather than providing an evaluation metric.
V-InFoR. [24] point out that “existing GNN explainers are not robust to the structurally corrupted graphs, namely graphs with noisy or adversarial edges.” This study investigates the negative effect of minor structural corruptions (which do not change the predictions) and severe structural corruptions (which change the predictions) on GNN explainers. Particularly, they provide quantitative evidence that existing GNN explainers are fragile to structurally corrupted graphs. They evaluate 6 GNN explainers under two settings: 1) minor corruptions where they select the perturbed graphs whose predictions are not changed after randomly flipping 20% edges; 2) severe corruptions where they employ adversarial attack algorithm GRABNEL [39] to perturb at most 10% edges to change the predictions. In contrast, our method focus on optimization-based (rather than random) and prediction-preserving (rather than prediction-altering) adversarial attacks.
RCExplainer. [16] aim to find a small set of edges of the input graph such that the prediction result will substantially change if we remove those edges. Specifically, they propose RCExplainer to generate robust counter-factual explanations for GNNs in two steps: 1) they employ a set of decision regions to model the decision logic of a GNN, where each decision region determines the predictions on multiple graphs that are predicted to be the same class; 2) they leverage a DNN model to explore the decision logic, and thereby extract robust counter-factual explanations as a small set of edges of the input graph. This method is only applicable to GNNs that belong to Piecewise Linear Neural Networks. In this paper, we focus on attacking differentiable post-hoc GNN explanation methods that can be applied to any GNNs.
GEAttack. [29] empirically demonstrate that GNN explanation methods can be employed as tools to detect the adversarial perturbations on graphs. On this basis, they propose an attack framework dubbed GEAttack that can jointly attack both GNN and its explanations. Specifically, they formulate GEAttack as a bi-level optimization problem: 1) they mimic the GNN explanation method optimization process to obtain graph explanations in the inner loop; 2) they compute the gradient of the attack objective w.r.t. the explanations in the outer loop. Different from GEAttack, we consider prediction-preserving attacks, which are inherently more challenging.
Concurrent Work. While the concurrent work [40] also addresses a similar topic, our study provides an alternative perspective by performing white-box attacks (namely we assume knowing the full knowledge about the GNN classifier and the GNN explainer), while they perform gray-box attacks by assuming knowing only the explanation loss and the generated explanatory edges of the GNN explainer.
II-D Adversarial Robustness of GNNs
Starting with the seminal works [41, 42], that both proposed adversarial attacks on GNNs, a rich literature formed in the realm of GNN-specific adversarial attacks, defenses, and certifications [43, 44]. Most attacks primarily focus on altering predictions via perturbations of the discrete graph structure (edge insertions/deletions), either between existing nodes [41] or via the insertion of new (adversarial) nodes [45]. Optimization-based attacks either operate globally (number of edge insertions/deletions) [46, 47], alter a local neighborhood [41], or can handle a combination of both constraints [48]. While adversarial attacks on GNN explanation methods may benefit from optimization methods over the discrete graph structure, a careful discussion and derivation of appropriate attack objectives and metrics was missing.
III Preliminaries
@@
III-A Node Classification with GNNs
We consider the task of node classification. Given an input graph $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathbf{X}\}$ (or $\mathcal{G}=(\mathbf{A},\mathbf{X})$), a prediction model $f_{\theta}(\cdot)$ assigns a label $c$ from a pre-defined set $\mathcal{C}$ to each node $v\in\mathcal{V}$. In other words, we have $f_{\theta}(\mathcal{G}(v))=c$, where $\mathcal{G}(v)$ denotes the computation graph of node $v$. Typically, $f_{\theta}(\cdot)$ consists of two components: 1) a GNN model used to learn node representations, and 2) a multi-layer perceptron (MLP) classifier that assigns labels. More specifically, given a set of training nodes and their labels $\{(v_{1},y_{1}),...,(v_{n},y_{n})\}$ from graph $\mathcal{G}$, the objective function of a GNN classifier can be defined as
@@ III-A Node Classification with GNNs
III-B GNN Explainer
We consider post-hoc GNN explanation methods that provide instance-level explanations: given a graph $\mathcal{G}$, a target node $v$, and prediction $f_{\theta}(\mathcal{G}(v))$, the explainer $g_{\lambda}(v,f(\mathcal{G}(v)))$ aims to ‘explain’ this particular prediction. Following common explanation methods, we assume that explanations are given in the form of an explanatory subgraph, namely $\mathcal{G}_{s}(v)=g(v,f(\mathcal{G}(v)))$, consisting of a small set of important edges and their associated nodes, where the edge importance is measured by a so-called edge importance score.
In this paper, we take PGExplainer [15] as an example to attack for three reasons. First, [49] pointed out that GNN explanations providing edge importance are preferable to explanations providing node importance, due to the fact that edges have more fine-grained information than nodes. Second, [50] empirically showed that PGExplainer generates the least unstable explanations, exhibiting a $35.35\%$ reduction in explanation instability relative to the average instability of other GNN explainers. Third, the objective function of PGExplainer is differentiable. For completeness, we briefly revisit the goal of PGExplainer. For brevity, we denote $\mathcal{G}(v)$ as $\mathcal{G}$ (and $\mathcal{G}_{s}(v)$ as $\mathcal{G}_{s}$ analogously) throughout the remainder of the paper.
@@ III-B GNN Explainer
IV GXAttack:
GNN Explanation Adversarial Attacks
Given an input graph $\mathcal{G}$, a GNN classifier $f_{\theta}(\cdot)$, and a GNN explainer $g_{\lambda}(\cdot)$, the original prediction for the target node $v$ is $f_{\theta}(\mathcal{G}(v))$ and its corresponding original explanatory subgraph is given by $\mathcal{G}_{s}(v)=g_{\lambda}(f_{\theta}(v,\mathcal{G}(v)))$. From an attacker’s perspective, a desirable manipulated graph $\mathcal{G}_{adv}:=\hat{\mathcal{G}}=\phi({\mathcal{G}})$ for attacking node $v$ should have the following properties, where $\phi(\cdot)$ is an attack function that perturbs its input, i.e., it removes and/or adds a small set of edges:
@@ IV GXAttack:
    The explanatory subgraph for node $v$ on original graph $\mathcal{G}_{s}(v)=g_{\lambda}(f_{\theta}(v,\mathcal{G}(v)))$ and that on perturbed graph $\hat{\mathcal{G}}_{s}(v)=g_{\lambda}(f_{\theta}(\hat{\mathcal{G}}(v)))$ are substantially different: $\mathcal{G}_{s}(v)\neq\hat{\mathcal{G}}_{s}(v)$. In particular, we aim to make $\hat{\mathbf{A}}_{s}$ (i.e., the adjacency matrix of explanatory subgraph on perturbed graph) strongly different from $\mathbf{A}_{s}$ (i.e., the adjacency matrix of explanatory subgraph on original graph).
IV-A Attack Objectives
@@
$\mathcal{L}:=~{}$ $\text{MI}(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\hat{\mathbf{A}},\mathbf{X}))$
$-\beta*\text{MI}(g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X})),g_{\lambda}(f_{\theta}(\hat{\mathbf{A}},\mathbf{X}))).$ (7)
@@
Final Objective. We directly model the edge perturbation matrix $\mathbf{P}\in\{0,1\}^{n\times n}$, where we flip $\mathbf{A}_{ij}$ (namely $0\to 1$ or $1\to 0$) if $\mathbf{P}_{ij}=1$, and keep $\mathbf{A}_{ij}$ intact otherwise. On this basis, Eq 6 can be rewritten as
$\underset{\mathbf{P}\in\{0,1\}^{n\times n}\text{~{}s.t.~{}}\sum\mathbf{P}_{ij}<B}{\max}~{}\text{MI}(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\mathbf{A}\oplus\mathbf{P},\mathbf{X}))$
$-\beta*\text{MI}(g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X})),g_{\lambda}(f_{\theta}(\mathbf{A}\oplus\mathbf{P},\mathbf{X}))),$ (8)
@@
IV-B Relaxations and Loss Definitions